Answer the question below based on the information given

. All data are in RM million

. Government expenditure function : 100

Investment function : 500

Consumption function : 200 + 0.6Yd

Tax function : 10 Export function : 0.2Y

Import function : 100

(i) Calculate the consumption and savings function after tax.

(ii) Calculate the national income equilibrium by using Injection = Leakage approach.

(iii)Calculate the national income equilibrium by using Aggregate Demand = Aggregate Supply.

(iv)Calculate the total consumption at the national income equilibrium.


1
Expert's answer
2022-01-29T11:04:41-0500

Solution:

i.). C = 200 + 0.6Yd = 200 + 0.6(Y – T) = 200 + 0.6Y – 6 = 194 + 0.6Y

Consumption function after tax = 194 + 0.6Y

S = Y – C

S = Y – (194 + 0.6Y)

S = Y – 194 – 0.6Y

S = -194 + Y(1 – 0.6)

S = -194 + 0.4Y

Saving function after tax = -194 + 0.4Y

 

ii.). The Injection = Leakage approach:

S + T + M = I + G + X

-194 + 0.4Y + 100 = 500 + 100 + 0.2Y

0.4Y – 0.2Y = 500 + 100 + 194 – 100

0.2Y = 694

Y = 3,470

Equilibrium national income = 3,470

 

iii.). AD = AS

Y = C + I + G + X – M

Y = 194 + 0.6Y + 500 + 100 + 0.2Y – 100

Y – 0.8Y = 694

0.2Y = 694

Y = 3,470

Equilibrium national income = 3,470

 

iv.). Total consumption at national income equilibrium:

C = 194 + 0.6Y

C = 194 + 0.6(3,470) = 194 + 2,082 = 2,276

Total consumption = 2,276